solve x-5/2 = x-1
(It reads x-5 over 2 = x-1)

Answer by Shana-D77(132)

Don't you just hate fractions?! So let's get rid of it:
(x-5)/2 = x-1
Multiply both sides by 2 to get rid of the /2:
x - 5 = 2(x - 1)
Distribute the 2(x - 1):
x - 5 = 2x - 2
-5 = x - 2 (subtracted x from both sides)
-3 = x (added 2 to both sides)
x = -3